# Stein Variational Evolution Strategies

### v1

@@ B.4 Reinforcement Learning
The goal of each RL task is to maximize the expected episodic return. The corresponding objective is to sample from the following posterior:
$p(\theta)\propto\exp(-J(\theta)),\quad J(\theta)=\operatorname*{\mathbb{E}}_{(s_{t},a_{t})\sim\pi_{\theta}}\bigl{[}\sum_{t=1}^{T}R(s_{t},a_{t})\bigr{]}.$
Here, we estimate the expected return via MC approximation across 16 independent policy rollouts for each iteration at which we evaluate the current particle set performance. In our experiments, we report the best expected return across all particles. Every such experiment is repeated 10 times, as with all previous tasks. We use the standard maximum episode length for all classic tasks as specified in the gym environment implementation, except for MountainCar, where we use a maximum episode length of 500 (instead of the default 1000 steps) to make the task more difficult. For the brax control tasks, we use a maximum episode length of 1000 steps. For the classic tasks Pendulum, CartPole and MountainCar, we use the gymnax implementations (Lange, 2022), and for the remaining environments, we use the brax-v1 implementations (Freeman et al., 2021). We use the continuous version of MountainCar. Please note that we use the brax legacy dynamics in our experiments.
@@ Appendix C Additional Results
C.1 Algorithm Ablations
Setting
[…]
Results
In Fig. 7, we display the results of our evaluation. When comparing SV-CMA-ES against single population CMA-ES we observe similar performances on most problems. The most notable difference that stands out is the better performance of SV-CMA-ES on the MountainCar problem. What happens in this experiment? The MountainCar problem is an MDP in which a car must be accelerated to reach a goal on a hill. Since there are negative rewards for large accelerations, there is a large local optimum at which policies are idle and do not accelerate at all to avoid negative rewards. However, once the goal is reached, the agent receives a reward of 100. We speculate due to these results that SV-CMA-ES may be superior to other CMA-ES-based methods on such sparse reward environments. We will further investigate this hypothesis in Appendix C.2. Furthermore, we compare the performance to multiple parallel CMA-ES runs. This experiment essentially ablates the kernel term in our update. We find that in comparison to multiple parallel CMA-ES runs, our method performs stronger on half of the problems, and equally well on the rest. These results are encouraging, as they suggest that our SV-ES is a viable alternative to independent parallelizations of a given ES algorithm.
@@ C.2 Environment Ablations
To investigate the hypothesis that SV-CMA-ES is superior to simple parallel CMA-ES runs or single large-population CMA-ES in sparse reward environments, we construct sparse reward versions of the Walker and Hopper environments following the procedure of Mazoure et al. (2019). The goal of the environment ablations is to make the optimization landscape more similar to that of MountainCar. To achieve this goal, we increase the control cost in the reward and remove the positive rewards for the agent being healthy. Additionally, the original forward reward is replaced by a positive reward for being upright that is only awarded once the agent has moved beyond a certain position in space. Essentially, the resulting environments only reward successful locomotion policies once the behavior is learned instead of rewarding intermediate learning outcomes as well. In this setting, the reward is negative for policies that move the agent not far enough and close to zero for policies that do not move the agent at all, similar to the MountainCar reward. The resulting reward function is:
@@
where $pos(s_{t})\in{\mathbb{R}}$ is the agent position at state $s_{t}$ in the task space. In our experiments, we choose $\alpha=2.0$ and $\beta=0.1$. Since the resulting ablated Walker problem is significantly more difficult, we run experiments on this benchmark for 1500 instead of 1000 iterations on this problem.
@@ Results
Our results show that SV-CMA-ES outperforms other CMA-ES-based methods considerably in the constructed scenarios. Both problems are more difficult to solve than the standard problems in Fig. 7, because a positive reward is only received once the agent moves far enough. As hypothesized, we observe a similar pattern as for MountainCar. In other words, SV-CMA-ES outperforms the other methods on sparse reward environments. These results suggest two things. First, we find that the coordination of multiple CMA-ES runs via the SVGD step improves the algorithmic performance on problems for which exploration is crucial. Second, we observe that this effect only comes to fruition if the population updates are coordinated via SVGD, as SV-CMA-ES outperforms parallel CMA-ES runs considerably. The combination of these findings suggests that SV-CMA-ES constitutes a strong alternative to other CMA-ES-based optimizers if problems require exploration. We thus believe that analysis of the scalability of the algorithm to a very large number of particles beyond our computational capabilities (i.e. several thousands) would be a very interesting direction of future research.

### v2

@@ B.4 Reinforcement Learning
The goal of each RL task is to maximize the expected return. The corresponding objective is to sample from the following posterior:
$p(\theta)\propto\exp(J(\theta)),\quad J(\theta)=\operatorname*{\mathbb{E}}_{(s_{t},a_{t})\sim\pi_{\theta}}\bigl{[}\sum_{t=1}^{T}r(s_{t},a_{t})\bigr{]}.$
Here, we estimate the expected return via MC approximation across 16 independent policy rollouts for each iteration at which we evaluate the current particle set performance. In our experiments, we report the best expected return across all particles. Every such experiment is repeated 10 times, as with all previous tasks. We use the standard maximum episode length for all classic tasks as specified in the gym environment implementation, except for MountainCar, where we use a maximum episode length of 500 (instead of the default 1000 steps) to make the task more difficult. For the brax control tasks, we use a maximum episode length of 1000 steps. For the classic tasks Pendulum, CartPole and MountainCar, we use the gymnax implementations [Lange, 2022], and for the remaining environments, we use the brax-v1 implementations [Freeman et al., 2021]. We use the continuous version of MountainCar. Please note that we use the brax legacy dynamics in our experiments.
@@ Appendix C Additional Results
C.1 Algorithm Ablations
Setting
[…]
Results
In Fig. 7, we display the results of our evaluation. When comparing SV-CMA-ES against single population CMA-ES we observe similar performances on most problems. The most notable difference that stands out is the better performance of SV-CMA-ES on the MountainCar problem. What happens in this experiment? The MountainCar problem is an MDP in which a car must be accelerated to reach a goal on a hill. Since there are negative rewards for large accelerations, there is a large local optimum at which policies are idle and do not accelerate at all to avoid negative rewards. However, once the goal is reached, the agent receives a reward of 100. We speculate due to these results that SV-CMA-ES may be superior to other CMA-ES-based methods on such sparse reward environments. We will further investigate this hypothesis in Appendix C.2.
Furthermore, we compare the performance to multiple parallel CMA-ES runs. This experiment essentially ablates the kernel term in our update. We find that in comparison to multiple parallel CMA-ES runs, our method achieves superior performance on half of the problems, and equal on the rest. These results are encouraging, as they suggest that our SV-ES is a viable alternative to independent parallelizations of a given ES algorithm.
Additionally, we include Fig. 8 to display the per-seed results on the MountainCar problem. The results highlight that GF-SVGD converges to the local optimum of being idle on two out of the ten runs.
Figure 8: Per-seed performances on the MountainCar task. SV-CMA-ES is the only method to achieve optimal performance across all seeds. GF-SVGD converges to idle agents (i.e., a reward of zero) on two out of the ten seeds.
@@ C.2 Environment Ablations
To investigate the hypothesis that SV-CMA-ES is superior to simple parallel CMA-ES runs or single large-population CMA-ES in sparse reward environments, we construct sparse reward versions of the Walker and Hopper environments following the procedure of Mazoure et al. [2019]. The goal of the environment ablations is to make the optimization landscape more similar to that of MountainCar. To achieve this goal, we increase the control cost in the reward and remove the positive rewards for the agent being healthy. Additionally, the original forward reward is replaced by a positive reward for being upright that is only awarded once the agent has moved beyond a certain position in space. Essentially, the resulting environments only reward successful locomotion policies once the behavior is learned, instead of rewarding intermediate learning outcomes as well. In this setting, the reward is negative for policies that move the agent not far enough and close to zero for policies that do not move the agent at all, similar to the MountainCar reward. The resulting reward function is:
@@
where $pos(s_{t})\in{\mathbb{R}}$ is the agent’s position at state $s_{t}$ in the task space. In our experiments, we choose $\alpha=2.0$ and $\beta=0.1$. Since the resulting ablated Walker problem is significantly more difficult, we run experiments on this benchmark for 1500 instead of 1000 iterations on this problem.
@@ Results
Our results demonstrate that SV-CMA-ES outperforms other CMA-ES-based methods considerably in the constructed scenarios. Both problems are more difficult to solve than the standard problems in Fig. 7 because a positive reward is only received once the agent moves far enough. As hypothesized, we observe a similar pattern as for MountainCar. In other words, SV-CMA-ES outperforms the other methods on sparse reward environments. These results suggest two things. First, we find that the coordination of multiple CMA-ES runs via the SVGD step improves the algorithmic performance on problems for which exploration is crucial. Second, we observe that this effect only comes to fruition if the population updates are coordinated via SVGD, as SV-CMA-ES outperforms parallel CMA-ES runs considerably. The combination of these findings suggests that SV-CMA-ES constitutes a strong alternative to other CMA-ES-based optimizers if problems require exploration. We thus believe that analysis of the scalability of the algorithm to a high number of particles beyond our computational capabilities (i.e., several thousands) would be an intriguing direction of future research.
@@
C.3 Empirical Runtime Analysis
Setting
Given the higher theoretical runtime complexity of SV-CMA-ES compared to the baselines, we perform an empirical runtime analysis. The goal of this investigation is to evaluate the significance of the theoretical algorithm properties in practice. To this end, we select multiple problems from the main paper to illustrate the performances over time. We select the synthetic Gaussian mixture as a low dimensional and fast-to-evaluate objective, a logistic regression tasks as moderately dimensional tasks with fast-to-evaluate objectives, and the brax RL tasks as the most high-dimensional task of the paper which require expensive simulator calls for evaluations of the ground truth density. For each problem, we run 1000 iterations, just as in prior experiments. For the RL tasks, we use the GPU accelerated brax simulator. To provide a fair analysis of wall-clock times, we disable deterministic operations on the GPU, which may lead to slight differences in performances, w.r.t. to the results reported in the main paper, but is more realistic in terms of runtimes. We would like to point out, however, that we did not observe any visible differences in the results. The experiment is carried out on a single NVIDIA A40 GPU and reported times are for all 10 independent runs being run in parallel.
Results
The results of the runtime analysis are depicted in Fig. 10. Overall, we observe competitive convergence behavior of SV-CMA-ES in terms of wall-clock time. While the overall runtime time required to run 1000 iterations is highest for SV-CMA-ES, we do observe that our method obtains good samples efficiently under this metric.
Generally, we observe that the influence of the update computation costs on the overall runtime varies greatly between problems. While the differences between all methods are relatively small on the synthetic 2d problems, we observe that SV-CMA-ES is slower on high dimensional problems such as the Covtype logistic regression task. This reflects that the computational cost of SV-CMA-ES increases with the number of problem dimensions, while the cost of the other baselines is dominated by the number of particles. However, we also observe that the quality of the updates of SV-CMA-ES is still generally better in these cases, as it samples well performing solutions the fastest.
More importantly, we observe that the main factor that drives overall runtime is the evaluation of the objective function. Furthermore, we observe smaller runtime differences between the methods on the Halfcheetah task despite it being the most high dimensional task that we evaluate on. This underlines that the significance of theoretical algorithm complexity depends heavily on the problem and the costs that are associated with target density evaluations. Since our method requires fewer iterations than the baselines to reach comparable performance (as shown previously), we find that this demonstrates the practical relevance of the presented approach.
(a) Gaussian Mixture
(b) Credit Log. Reg.
(c) Covtype Log. Reg.
(d) Halfcheetah RL
(e) Hopper RL
(f) Walker RL
Figure 10: Performance vs. wallclock time. We run all methods for 1000 iterations and display the elapsed wallclock time. The plot shows that SV-CMA-ES also performs well w.r.t. this metric. All results are averaged across 10 independent runs (±$\pm 1.96$ standard error).
(a) Gaussian Mixture
(b) Double Banana
(c) Motion Planning
Figure 11: Mean step lengths per iteration. We record the step length every 20 steps across 1 00010001\,0001 000 iterations. The SV-CMA-ES steps and estimated gradient steps by CMA-ES converge to a stable equilibrium. We report the mean (±$\pm 1.96$ standard error) across all 100 particles and 10 independent runs.
C.4 Empirical Convergence Analysis
To evaluate the convergence of our method, we expand our analysis beyond the pure MMD computations in Fig. 3 and analyze the length of the particle steps. To this end, we use the same setting as the one in Fig. 3 to plot the Frobenius norm of the particle preconditioning matrix $\sigma C^{1/2}$ across iterations in Fig. 11. In the same figure, we further plot the length of the steps that were actually taken by the particles, i.e., the size of the update in Eq. (15). The results show that the algorithm exhibits stable convergence. A key finding in this context is that for each problem, there exists a stationary point at which the CMA-ES steps that are sampled from ${\cal N}(\mathbf{x},\sigma^{2}C)$ are counteracted by the kernel gradients, as the total step length is shorter than the matrix norm. This shows that our method is capable of finding the point of convergence of SVGD where the kernel gradient and objective gradient balance each other out. We see the fact that the resulting total step length does not reach zero across all problems as an artifact of the stochastic gradient estimates that CMA-ES provides. We do not see this as a problem, however, as the MMD analysis in Fig. 3 shows that the particles only move within areas of high density. Since the approximation of the target density in SVGD is based on a sum of delta function which approximates the target as $q(x)=\sum_{\mathbf{x}_{i}\in X}\delta(\mathbf{x}-\mathbf{x}_{i})/\varrho$ it permits small particle movements at the equilibrium between repulsive and driving force in the update.
C.5 Convergence Plots
For completeness, we depict the convergence results on the synthetic sampling tasks. For each task and method combination, we display the evolution of the sample set across the optimization. To this end, we split the total number of iterations into 10 equally-sized bins and generate a plot for each. In other words, we plot the sample set every 100 iterations. The results for the Gaussian mixture are listed in Fig. 12, those for the Double Banana density in Fig. 13, and Fig. 14 shows the convergence on the trajectory optimization task. All plots confirm the quantitative results, i.e., SV-CMA-ES converges quickly compared to the baselines. We note that we refer to convergence once all particles are in high density areas. Due to the stochastic nature of the update, there is no guarantee that there is an actual stationary distribution of the finite optimization process. For storage reasons, we include these images as raster images. We encourage readers to reach out to us in case they are interested in the vector graphics version for these last three plots.
(a) $\nabla$-SVGD
(b) SV-CMA-ES
(c) GF-SVGD
(d) SV-OpenAI-ES
Iteration 15001000
Figure 12: Convergences on GMM sampling task. For each method, the sample convergence across the full 1000 sampling iterations is displayed.
(a) $\nabla$-SVGD
(b) SV-CMA-ES
(c) GF-SVGD
(d) SV-OpenAI-ES
Iteration 15001000
Figure 13: Convergences on Double banana sampling task. For each method, the sample convergence across the full 1000 sampling iterations is displayed.
(a) $\nabla$-SVGD
(b) SV-CMA-ES
(c) GF-SVGD
(d) SV-OpenAI-ES
Iteration 15001000
Figure 14: Convergences on motion planning sampling task. For each method, the sample convergence across the full 1000 sampling iterations is displayed.